# Skill vs. Chance Quantification for Popular Card & Board Games

### v1

@@ Abstract
We consider a few online and offline games under actual playing conditions. Generally it is expected that initially a player obtains additional skill with experience of playing more number of games and then it should finally saturate. This phase is identified when a player, with the experience of very few games, loses more when she plays against players with much longer history. Then the winning proportion curve moves up and finally it saturates. We benchmark our analysis and discussion against Chess, the most skilled one among the games we consider here. We use proprietary data from actual games (online and offline) as well as experiments for our statistical analysis. In this regard, we show that Rummy has stronger skill and learning effects. Ludo has similar characteristics as Rummy, but at a weaker level. Similarly, a game that is perceived as almost no skill such as Teen Patti indeed presents much less skill in the analysis. In the next section we describe the game structures.
@@ 1 Introduction
In this world, games serve as both entertainment and cognitive exercises, fostering strategic thinking and social interaction. Whether in the online or offline mode, they offer a diverse spectrum of experiences, engaging players in skillful pursuits and immersive narratives. Today, games are also a medium of business, drawing players from all corners of the world and impacting the economy (see [1] for survey based overview for the online Rummy game). Owing to the drastic reduction in costs of smart phones and internet usage and driven by the expansion of several interactive digital platforms, a huge market has grown for the online games among teenagers and adults.
Certain factors involved in a game could be evaluated from the player level data of the past games to resolve this question. Online games, in an Indian context, may belong to the category of the games of skill provided the success (measured, for example, by the winning rate of the same player) in such games depends on training and learning ability through experience and prowess of the players (A large literature has developed on this over the last few years, see for example [2], [3] and [4]). More specifically see [5] and [6] for analysis of digital games.
We conduct an academic study to assess the materiality of “skill” based on real user data in a few popular online card and board games. This draft analyzes the ”skill” factor associated with these games based on the winning statistics of a set of players who play against each other and have different levels of experience and capabilities [7, 3]. We take a frequentist’s approach as in [8] and showcase several elaborate results on the data gathered from a major online platform that host these games. Our Luck vs. Law modelling methodology is described in Section 3. The four games under consideration are statistically analysed in the subsequent sections sequentially. Finally, a comparative discussion is made in section 8.
2 The Game Structures Considered
@@
The quantification of skill vs. chance in a game has several inputs. We try to arrange these in a logical manner below.
@@
    Randomisation only at the beginning: we start with the unlikely example of Chess, which is considered to be a purely skill based game. The only role of chance is playing with Black or White pieces, which is statistically shown to have some implication in terms of winning ceteris paribas (discussed earlier).
    A second, and more common example is that of Rummy where the role of chance is in the initial permutation of the cards. Then a partition of which results in the two hands with the players and the sequence in the deck. We have already shown that there is significant skill (experience) premium in the Rummy game.
@@
    There is randomness in game play also: Ludo is a prime example where the players start with identical initial positions but nature’s move (roll of the dice) dictates the choices available to them at each turn. Again, we have shown above that substantial skill premium (benefit from the use of smart strategy) exists in Ludo.
@@
    A second aspect is observability: whether the players can see opponent’s position fully (as in Chess or Ludo) or not (in Rummy). This results into mental probability/belief updations that are refined through the course of game play. this is another aspect of Rummy that requires substantial skill.
@@
    Thus one may think of variations on the existing games which would highlight one or the other aspects (randomisation, observability) of a game more. Consider the Ludo Ninja variation (Zupee) which decides the rolls of the die for each turn at one go. At the very beginning a sequence of numbers in $\{1,2,3,4,5,6\}$ are given to each player. So there is no randomisation after the start. The sequence for the opponent is unknown to a player. So there is also incomplete observability. In this sense, Ludo Ninja is closer to Rummy. This variation of Ludo also benefits from the use of smart strategy.
@@
    One may also think of a further variation where both the lists will be observable to the players, so that there is no lack of observability and then the game becomes one of pure strategy like Chess. Experimentation is required to understand the value of skill in this version.
So how do we proceed towards this quantification? We adopt an approach similar, but not identical, to the Getty et. al. (2018) paper in the context of fantasy games or that of [9] for Poker. Our approach will be more statistically oriented, through a regression type analysis. In this, we follow more closely our previous study on Online Rummy Game, [8]. For this purpose it is imperative to create a benchmark for Chess, which is considered closest to a pure skill game. We will use statistical methodology similar to that used for Rummy in the context of the game of Chess where skill will be proxied by ELO rating and age (in a non-monotonic manner, see [10]).
@@
Data: For Rummy, Ludo and Teen Patti, we use proprietary data from OSG companies. For benchmarking purpose we also do extensive simulation studies. For Chess, we use publicly available game level outcome data from FIDE official website.
3 Luck vs. Law measurement formulation
@@
  • Learning $l_{t}=l(e_{t})$, where $e_{t}$ is experience, observable.
    Learning rate can be different between players (efficiency). But also need to consider fatigue.
@@
Estimation Strategy: As skill has unobservable inputs, a direct estimation of the above relationship from (experience, outcome) data is not possible. So we posit the following:
$dw_{t}=\frac{\partial w}{\partial h}+\frac{\partial w}{\partial l_{t}}$ and $dw_{t-1}=\frac{\partial w}{\partial h}+\frac{\partial w}{\partial l_{t-1}}.$ Differencing obtains $dw_{t}-dw_{t-1}=\frac{\partial w}{\partial l_{t}}-\frac{\partial w}{\partial l_{t-1}}$
As we are looking for an intuitive and easily computable benchmarking procedure, we stick to a linerised relationship. Hence, we posit the following regression equation to represent a linear approximation of the above DE:
$w_{t}=\alpha+\beta_{1}w_{t-1}+\beta_{2}e_{t-1}+\beta_{3}\Delta e_{t}+\text{error}$
(here experience works as a proxy for learning).
First, consider Rummy, Ludo and TeenPatti. Consider two time intervals (0, 1) and (1, 2). suppose experience at these time points are $m_{i},\,i=0,1,2.$ Then calculate $e_{1}=(m_{0}+m_{1})/2$ and $e_{2}=(m_{1}+m_{2})/2-m_{1}=(m_{2}-m_{1})/2.$ Also consider the win percentage in interval (0, 1) is $w_{1}$ and in (1, 2) is $w_{2}.$ For empirical convenience, we will use the following transformed version of the regression equation, that spans the whole real line:
$\Phi^{-1}(w_{2})=\alpha_{0}+\alpha_{1}\Phi^{-1}(w_{1})+\beta_{1}e_{1}+\beta_{2}e_{2}+\text{error}$
For Chess, we will use FIDE rating at the start of the two tournaments as $e_{1}$ and $e_{2}$.
4 Analysing Chess Tournament Data
@@ 4.1 Classical
For Classical format Chess, the leading FIDE tournament data is used. The downside of this being very small number of common players (the tournament only allows the top 14 players, see table 1). Thus results are not very reliable. In fact, none of the explanatory factors turn out to be significant (see table 2). So we continue to a version with a larger pool of common players.
Player 1st Year Win % 2nd Year Win % $\phi^{-1}$(1st Year Win %) $\phi^{-1}$(2nd Year Win %) 1st Year Rating 2nd Year Rating

### v2

# Skill vs. Chance Quantification and Relative Ranking for Popular Card & Board Games

@@ Abstract
This paper presents a data-driven statistical framework to quantify the role of skill in games, addressing the long-standing question of whether success in a game is predominantly driven by skill or chance. We analyze player-level data from four popular games—Chess, Rummy, Ludo, and Teen Patti—using empirical win statistics across varying levels of experience. By modeling win rate as a function of experience through a regression framework and employing empirical bootstrap resampling, we estimate the degree to which outcomes improve with repeated play. To summarize these dynamics, we propose a flexible skill score that emphasizes learning over initial performance, aligning with practical and regulatory interpretations of skill. Our results reveal a clear ranking, with Chess showing the highest skill component and Teen Patti the lowest, while Rummy and Ludo fall in between. The proposed framework is transparent, reproducible, and adaptable to other game formats and outcome metrics, offering potential applications in legal classification, game design, and player performance analysis.
Keywords: Chance, Chess, Ludo, Rummy, Skill, Statistical Analysis, Teen Patti
@@ 1 Introduction
Games, whether played online or offline, serve not only as entertainment but also as exercises in strategy, engaging players in decision-making, learning, and social interaction. The rapid growth of the online gaming industry driven by the ubiquity of smartphones and internet access has transformed games into a multi-billion-dollar business ecosystem. However, this expansion has also raised legal, economic, and ethical concerns, particularly when monetary rewards are involved. Ethical considerations in the games industry are gaining prominence, with ongoing discourse emphasizing the need for responsible design, player well-being, and inclusive development practices [1].
At the heart of legal and regulatory debates is the classification of games as either skill-based or chance-based, a distinction with far-reaching policy and financial implications. Game mechanics such as levels and high scores have been shown to significantly influence user engagement and persistence, highlighting the motivational power of structured progression systems within gamified experiences [2]. Yet, the boundary between skill and chance is often blurred, as most games incorporate elements of both. Thus, the critical question is not whether a game is purely one or the other, but rather: how much skill is involved in determining outcomes? Quantifying this continuum with empirical rigor is essential for informed debate, legal clarity, and greater industry transparency.
Several recent works have addressed this challenge using diverse approaches. Goodman et al. (2024) propose a simulation-driven framework to separate skill and randomness in tabletop games, emphasizing how tournament seeding impacts measurable outcomes [3]. Duersch et al. (2020) provide an econometric variance decomposition model to quantify the relative importance of skill and chance, especially for regulatory classification of games [4]. Borm and van der Genugten (2001) introduce a relative measure of skill for games with chance elements by comparing expected and observed success frequencies across player types [5]. Getty et al. (2018) present a method to quantify the legal implications of randomness in fantasy sports using information-theoretic concepts [6]. Levitt and Miles (2014) analyze extensive poker data to statistically demonstrate the presence of skill over repeated play [7]. Orkin (2021) explores the philosophical and mathematical boundaries between skill and chance, highlighting the role of statistical repeatability as an indicator of skill [8]. Several additional works have explored the complex boundary between skill and chance in games through various empirical and theoretical lenses. Agt et al. (2021) present a student-led empirical analysis comparing different games on a skill-chance spectrum using win/loss data and structural game features [9] . Erdmann (2011) provides a theoretical framework for characterizing games based on the degree of player influence versus randomness, emphasizing philosophical and probabilistic perspectives [10]. Hannum and Cabot (2009) argue for the legal classification of poker as a game of skill, synthesizing statistical evidence with legal criteria [11]. More recently, Stafford and Vaci (2022) discuss how digital games can be leveraged to study skill acquisition over time, advocating for fine-grained behavioral data in psychological research [12]. These works underscore the diversity of approaches to measuring skill and support the need for a scalable, data-driven framework like the one proposed in this paper.
Building on this literature, our paper presents a statistically grounded, data-driven framework to quantify the role of skill in popular card and board games using real-world player-level data. Specifically, we analyze four games: Chess, Rummy, Ludo, and Teen Patti , to determine the extent to which experience and prior performance influence success. We define ‘skill’ as a player’s ability to consistently improve outcomes through learning, adaptation, and strategy, in contrast to outcomes primarily driven by randomness or fixed initial conditions.
Our contribution is threefold. Firstly, we propose a regression-based skill quantification model that uses transformed win rates and player experience as key inputs, accounting for both innate ability and learning. Secondly, we introduce an empirical bootstrap resampling procedure to assess the robustness and uncertainty of skill estimates across games, avoiding reliance on strong parametric assumptions. Finally, we construct a flexible skill score metric that allows for subjective weightings and cut-offs, enabling comparative ranking of games while maintaining transparency about the assumptions involved.
The framework is applied to large-scale proprietary datasets (Rummy, Ludo, Teen Patti) and public datasets (Chess) to derive skill scores for each game. Our results show that while Chess predictably scores highest on the skill spectrum, games like Rummy and Ludo also demonstrate significant skill components—contrary to popular perception—whereas Teen Patti shows relatively weak evidence of strategic learning.
The remainder of this paper is structured as follows. Section II presents the algorithmic framework in a step-by-step format to enhance transparency and reproducibility. Section III outlines the empirical methodology, including the model specification, estimation strategy, and resampling techniques used to derive robust skill estimates. In Sections IV through VII, we present game-wise statistical analyses of Chess, Rummy, Ludo, and Teen Patti, respectively—detailing both regression and bootstrap results. Section VIII introduces our flexible scoring methodology to convert statistical findings into interpretable skill scores and provides comparative rankings. Finally, Section IX discusses key insights, limitations, and potential extensions, including the applicability of the framework to alternative measures of game success.
1.1 The Game Structures Considered
@@
The quantification of skill vs. chance in a game has several inputs. We arrange these in a logical manner as below.
@@
    Randomisation only at the beginning: we start with the unlikely example of Chess, which is considered to be a purely skill based game. The only role of chance is playing with Black or White pieces, which is statistically shown to have some implication in terms of winning ceteris paribas (a Latin phrase that means “all other things being equal" or “other things held constant").
    A second, and more common example is that of Rummy, where the role of chance is in the initial permutation of the cards. Then a partition of which results in the two hands with the players and the sequence in the deck. Here, it is shown that there is significant skill (experience) premium in the Rummy game.
@@
    There is randomness in game play also: Ludo is a prime example where the players start with identical initial positions but nature’s move (roll of the dice) dictates the choices available to them at each turn. Again, we show that substantial skill premium (benefit from the use of smart strategy) exists in Ludo.
@@
    A second aspect is observability: whether the players can see opponent’s position fully (as in Chess or Ludo) or not (in Rummy). This results in mental probability/belief updations that are refined through the course of game play. this is another aspect of Rummy that requires substantial skill.
@@
    Thus one may think of variations on the existing games which would highlight one or the other aspects (randomisation, observability) of a game more. Consider the Ludo Ninja variation (https://www.zupee.com/ludo/ludo-ninja/) which decides the rolls of the die for each turn at one go. At the very beginning, a sequence of numbers in $\{1,2,3,4,5,6\}$ are given to each player. So there is no randomisation after the start. The sequence for the opponent is unknown to a player. So there is also incomplete observability. In this sense, Ludo Ninja is closer to Rummy. This variation of Ludo also benefits from the use of smart strategies.
@@
    One may also think of a further variation where both the lists will be observable to the players, so that there is no lack of observability and then the game becomes one of pure strategy like Chess. Experimentation is required to understand the value of skill in such versions.
Thus, how do we proceed towards this quantification? We adopt an approach similar, but not identical, to [6] in the context of fantasy games or that of [7] for Poker. Our approach will be more statistically oriented, through a regression type analysis. For this purpose it is imperative to create a benchmark for Chess, which is considered closest to a pure skill game. We will use statistical methodology similar to that used for Rummy in the context of the game of Chess where skill will be proxied by ELO rating and age (in a non-monotonic manner, see [13]). The advantages of this approach is threefold. First, the change dynamics of performance is captured in terms of only a few parameters (see equation (1) below) like experience as proxy for learning, previous performance as proxy for innate skill and current performance to model saturation (fatigue). Secondly, we use a linearised version of this for empirical purposes that enables the use of the easily understood linear regression methodology (see equation (2) below) that allows the study of partial effects easily. Thirdly, our skill quantification formulation allows sufficient flexibility in choosing the sensitivity with respect to each of these parameters in terms of scaling and critical cut-offs (see eqn (3) below) as well as their relative importance (see equation (4) below). This flexibility results in a distribution of skill scores rather than a single number, as evidenced in table 21 and figures 6, 7, 8 and 9. Is this unsatisfactory? We claim that this is more useful than a single number as opinions vary widely in terms of relative importance and scaling of these parameters in the skill measurement literature. Our framework provides, to the best of our knowledge, the first objective data driven framework to pursue this debate in a scientific manner.
@@
Data: For Rummy, Ludo and Teen Patti, we use proprietary data from OSG companies (in particular Games 24X7, https://www.games24x7.com/). For benchmarking purpose we also do extensive simulation studies. For Chess, we use publicly available game level outcome data from FIDE official website (https://www.fide.com/).
2 Algorithmic Approach
To enhance transparency and reproducibility, we outline below the step-by-step procedure followed in our analysis. This algorithmic presentation is designed to be easily understood, even by readers without a technical background.
Step 1: Data Preparation
  • Collect player-level data for each game, including the number of games played and corresponding win rates.
  • Organize data into intervals based on the number of games played (e.g., 1–10 games, 11–20 games, etc.).
Step 2: Experience Grouping
  • Divide players into experience quantiles according to the number of games played.
  • For each quantile, compute the outcome (e.g., average win rate) to track how performance evolves with experience.
  • Samples are required to be created at a user level across different experience partitions and/or within partitions to create enough variety for the modeling purposes
  • One manifestation of this could be to consider the experience (i.e., the no. of games) into two halves and try to infer the second half outcome as a function of the first one. This can be performed across different experience levels to ensure the variety (e.g., players playing 100 games being partitioned into 1st 50 and 2nd 50, players playing 500 games being partitioned into 1st 250 and 2nd 250, and so on).
Step 3: Regression Modeling
  • Fit a regression model where the dependent variable is the win rate and the independent variable is the experience level. (The model is at a user level. We are using the groupings in the third point of step 2. Partitions refer to the two halves based on which player performance and experience are considered. )
  • Estimate the slope to quantify the skill-growth trend - how much a player’s win rate improves with additional experience.
Step 4: Bootstrap Resampling
The buckets refer to the groupings in the third point of step 2. Resampling is done over the set of all users as we have performance and experience data of two halves for all users.
  • Apply the empirical bootstrap technique within each experience bucket to estimate the uncertainty around the average win rate.
  • For each resample, recalculate the average win rate to generate a distribution of estimates.
Step 5: Confidence Interval Construction
Confidence intervals are needed to provide the estimate of probability perspective (uncertainty, fluctuations etc.) of a model.
  • Use bootstrap distributions of regression slopes to construct confidence intervals. (Construction of confidence intervals using bootstrap is preferred since it doesn’t require any parametric assumptions.)
  • These intervals quantify the statistical uncertainty associated with the estimated skill-growth rate.
Step 6: Normalize Skill Estimates
For comparability across games and avoid possible scaling issues, a data dependent normalisation of the estimates is necessary.
  • For each game, normalize the bootstrap slope estimates by dividing them by the average win rate in the first experience half.
  • This yields a relative skill slope that accounts for baseline win rate differences across games.
Step 7: Compute Summary Score
  • The median of the outcome measure (e.g., win rate) can be considered as a summary, as it is more robust to outliers than the conventional average value. The model was run a large number of times on resampled datasets, and median here refers to the median of the summary statistics from the repeated executions of the model.
  • This median represents the Final Skill Score—a single interpretable value capturing the game’s skill component.
  • Compute the median of the relative slope distribution for each factor for each game.
Step 8: Cross-Game Comparison
  • Use the Final Skill Scores to rank or compare games in terms of how much skill contributes to success.
  • Higher scores suggest greater influence of skill, while lower scores may imply chance, volatility, or short learning curves.
3 Framework for Empirical Analysis
3.1 Data Format
To begin with consider having a games outcome data, measured in terms of win/loss for various games for multiple players. For an empirical investigation of the robustness of our findings, we need a tournament setting where a large pool of players play against each other. These results can be used to empirically define the variability (or otherwise) of our results using resampling techniques.
3.2 Our Luck vs. Law measurement formulation
After suitably identifying the data, a suitable statistical model is needed to determine whether the game is of skill or chance.
@@ 3.1 Data Format
    • Learning $l_{t}=l(e_{t})$, where $e_{t}$ is experience, measured by the number of games previously played and so it is observable.
    Learning rate can be different between players (efficiency). However, one needs to consider the fatigue too.
@@
3.3 Estimation Strategy
As skill has unobservable inputs, a direct estimation of the above relationship is not possible from (experience, outcome) data. So we posit the following:
$dw_{t}=\frac{\partial w}{\partial h}+\frac{\partial w}{\partial l_{t}}$ and $dw_{t-1}=\frac{\partial w}{\partial h}+\frac{\partial w}{\partial l_{t-1}}.$ Differencing yields $dw_{t}-dw_{t-1}=\frac{\partial w}{\partial l_{t}}-\frac{\partial w}{\partial l_{t-1}}$. Denote $\Delta e_{t}=e_{t}-e_{t-1}$.
As we are looking for an intuitive and easily computable benchmarking procedure, we stick to a linearized relationship. Hence, we posit the following regression equation to represent a linear approximation of the above Differential Equation:
$w_{t}=\alpha+\beta_{1}w_{t-1}+\beta_{2}e_{t-1}+\beta_{3}\Delta e_{t}+\text{error}$ (1)
(here experience works as a proxy for learning and the outcome represents the chance).
Consider two time intervals (0, 1) and (1, 2). Denote by $m_{i},\,i=0,1,2$ experience (games previously played) at these time points. Calculate $e_{1}=(m_{0}+m_{1})/2$ and $e_{2}=(m_{1}+m_{2})/2-m_{1}=(m_{2}-m_{1})/2$. Denote the win percentage in intervals (0,1) and (1,2) as $w_{1}$ and $w_{2}$.
Practical Application: For empirical convenience, we will use the following transformed version of the regression equation, that spans the whole real line:
$\Phi^{-1}(w_{2})=\alpha_{0}+\alpha_{1}\Phi^{-1}(w_{1})+\beta_{1}e_{1}+\beta_{2}e_{2}+\text{error}.$ (2)
, where $\Phi^{-1}(w)$ is the represents the inverse of the standard normal cumulative distribution function, also known as the quantile function.
3.4 Assessing Robustness using Resampling
We employed the empirical bootstrap technique to address the regression problem, enabling robust estimation of variability without relying on strong parametric assumptions. In this case, the empirical bootstrap is also called paired bootstrap. Consider given a paired sample: $\left(X_{1},Y_{1}\right),\cdots,\left(X_{n},Y_{n}\right)$, (where $X_{i},Y_{i}$ represent some data points) we generate new sets of IID observations
$\left(X_{1}^{*},Y_{1}^{*}\right),\cdots,\left(X_{n}^{*},Y_{n}^{*}\right)$
such that for each $\ell$,
$P\left(X_{\ell}^{*}=X_{i},Y_{\ell}^{*}=Y_{i}\right)=\frac{1}{n},\quad\forall i=1,\cdots,n.$
Namely, we treat $\left(X_{i},Y_{i}\right)$ as one object and sample with replacement $n$ times from these $n$ objects to form a new bootstrap sample. Thus, each time we generate a set of $n$ new observations from the original dataset. Assuming the entire process is repeated $B$ times, we would obtain
$\left(X_{1}^{*(1)},Y_{1}^{*(1)}\right),\cdots,\left(X_{n}^{*(1)},Y_{n}^{*(1)}\right)\\ \left(X_{1}^{*(2)},Y_{1}^{*(2)}\right),\cdots,$
$\left(X_{n}^{*(2)},Y_{n}^{*(2)}\right)\left(X_{1}^{*(B)},Y_{1}^{*(B)}\right),\cdots,\left(X_{n}^{*(B)},Y_{n}^{*(B)}\right)$
$B$ sets of fitted coefficients. We then estimate the variance by
$\widehat{\operatorname{Var}}_{B}\left(\widehat{\beta}_{0}\right)=\frac{1}{B}\sum_{\ell=1}^{B}\left(\widehat{\beta}_{0}^{*(\ell)}-\bar{\beta}_{0}^{*}\right)^{2},\quad\bar{\beta}_{0}^{*}=\frac{1}{B}\sum_{\ell=1}^{B}\widehat{\beta}_{0}^{*(\ell)}$ ,$\widehat{\operatorname{Var}}_{B}\left(\widehat{\beta}_{1}\right)=\frac{1}{B}\sum_{\ell=1}^{B}\left(\widehat{\beta}_{1}^{*(\ell)}-\bar{\beta}_{1}^{*}\right)^{2},$ $\quad\bar{\beta}_{1}^{*}=\frac{1}{B}\sum_{\ell=1}^{B}\widehat{\beta}_{1}^{*(\ell)}$.
[…]
In summary, the application of the empirical bootstrap framework allows us to reliably estimate the variability in skill indicators across players with differing experience levels, without relying on strong distributional assumptions. This lays a statistically sound foundation for comparing skill dynamics across different games in the subsequent sections
4 Analyzing Chess
Chess is a two-player, turn-based strategy game played on an $8\times 8$ board. It is a game of perfect information, meaning all elements of the game state are visible to both players at all times. The objective is to checkmate the opponent’s king. With no inherent randomness and a wide decision space, Chess is widely recognized as a skill-intensive game requiring deep planning and pattern recognition.
We present the analysis for various cases as follows.
@@ 4.1 Classical
For Classical format Chess, the leading FIDE tournament data is used. The downside of this being very small number of common players (the tournament only allows the top 14 players, see table 1). Thus results are not very reliable. In fact, none of the explanatory factors turn out to be significant as the p-value is greater than $5\%$ for all the covariates(see table 2). So we continue to a version with a larger pool of common players.
Player 1st Year Win % ($w_{1}$) 2nd Year Win % ($w_{2}$) $\phi^{-1}(w_{1})$ $\phi^{-1}(w_{2})$ 1st Year Rating ($e_{1}$) 2nd Year Rating ($e_{2}$)